# The Empty Chair: Using LLMs to Raise Missing Perspectives in Policy Deliberations

### v1

@@ 7. Discussion
In this work, we explored the potential and risks of using AI-generated personas to introduce missing perspectives into small-group policy deliberations. Overall, participants responded positively, finding the tool useful for expanding their consideration of absent viewpoints and prompting discussions in new directions. Many appreciated how the personas made abstract concerns more concrete and surfaced overlooked perspectives, with facilitators noting that the tool encouraged deeper, more candid engagement. Delegates and facilitators both suggested that having the personas more dynamically respond to the conversation would be an interesting and useful addition to the current tool. However, two key concerns emerged. First, participants worried that AI-generated personas could misrepresent certain groups, potentially replacing rather than complementing authentic voices. Second, some found the personas’ contributions too generic or redundant, limiting their impact. Facilitators suggested that explicitly framing the tool either as a role-playing aid or as a mechanism for surfacing overlooked perspectives could help clarify its purpose. Future iterations could improve persona specificity and grounding by incorporating real-world data, such as community input or structured datasets, to enhance authenticity and relevance. While AI has the potential to enrich deliberation by introducing different viewpoints, careful design is needed to ensure it serves as a supplement rather than a substitute for genuine representation. We hope this work provides useful insights into the benefits and risks of deploying AI personas in a realistic deliberative settings and helps spark ideas for new applications of LLMS and personas in deliberations.
8. Limitations
@@ 9. Ethical Considerations
Acknowledgements.
Appendix A Technical Details
For all LLM responses, we used GPT-4o. Given that LLMs may have different capabilities in modeling various perspectives (Santurkar et al., 2023) we hope that in the future tools utilizing AI generated personas have various base model options that users can choose. Our app was deployed on an AWS EC2 instance that could be accessed from a public URL by facilitators.
Appendix B Prompts
@@
B.1. Stakeholder Generation Prompt
@@
The stakeholders should be individuals who may have a vested interest in the university’s sustainability policies but are not undergraduate students. For example, they may be other types of university affiliates, people in the Boston area, or individuals who are not affiliated with the university but are still impacted by sustainability policies. At least one stakeholder should have low sustainability interest. Format the response as a JSON object with a ’stakeholders’ array containing objects with ‘‘name’’, ‘‘description’’, and ‘‘demographics’’ fields. The conversation is as follows:
@@
B.2. Stakeholder Reflections Prompt
@@
Appendix C Generating Stakeholder Questions

### v2

@@ 7 Discussion
In this work, we explored the potential and risks of using AI-generated personas to introduce missing perspectives into small-group policy deliberations. Overall, participants responded positively to AI generated personas, finding the tool useful for expanding their consideration of absent viewpoints and prompting discussions in new directions. Many appreciated how the personas made abstract concerns more concrete and surfaced overlooked perspectives, with facilitators noting that the tool encouraged deeper, more candid engagement. Delegates and facilitators both suggested that having the personas more dynamically respond to the conversation would be an interesting and useful addition to the current tool. However, two key concerns emerged. First, participants worried that AI-generated personas could misrepresent certain groups, potentially replacing rather than complementing authentic voices. Second, some found the personas’ contributions too generic or redundant, limiting their impact. Facilitators suggested that explicitly framing the tool either as a role-playing aid or as a mechanism for surfacing overlooked perspectives could help clarify its purpose and reduce the chance it would be viewed as replacing participants. Future iterations of this approach could improve persona specificity and grounding by incorporating real-world data, such as community input or structured datasets, to enhance authenticity and relevance. While AI has the potential to enrich deliberation by introducing different viewpoints, careful design is needed to ensure it serves as a supplement rather than a substitute for genuine representation. We hope this work provides useful insights into the benefits and risks of deploying AI personas in a realistic deliberative settings and helps spark ideas for new applications of LLMs and personas in deliberations.
Appendix A Limitations
@@ Appendix B Ethical Considerations
Appendix C Additional Figures
We have added a flow diagram of the context our tool received and what outputs it generated (see Figure 2). We also show the reflection of the AI persona (see Figure 3) and an example of a generated question (see Figure 4). Finally, we plot the average results for our survey responses in Figure 6 and Figure 5.
Figure 2: The flow for generating missing stakeholders, stakeholder reflections, and finally stakeholder questions based on the assembly context and a live transcription of the conversation. The solid lines capture how the assembly context and the transcript helps generate missing stakeholder personas, which then is used to generate stakeholder reflections, and then finally the stakeholder questions.
Figure 3: An example of points of disagreement and missing perspectives from a persona.
Figure 4: An example question generated by a persona.
Figure 5: Changes in understanding towards different points between pre-survey (prior to first day) and post-survey (after first day). Red lines capture pre/post means, and blue lines and dots represent individuals.
Figure 6: Survey responses capturing how participants felt about various aspects of the activity as well as their overall impressions
Appendix D Technical Details
For all LLM responses, we used GPT-4o. Given that LLMs may have different capabilities in modeling various perspectives santurkar2023opinionslanguagemodelsreflect we hope that in the future tools utilizing AI generated personas have a variety of base model options from which users can choose. Our app was deployed on an AWS EC2 instance that could be accessed from a public URL by facilitators.
Appendix E Prompts
@@
E.1 Stakeholder Generation Prompt
@@
The stakeholders should be individuals who may have a vested interest in the university’s sustainability policies but are not undergraduate students. For example, they may be other types of university affiliates, people in the area, or individuals who are not affiliated with the university but are still impacted by sustainability policies. At least one stakeholder should have low sustainability interest. Format the response as a JSON object with a ’stakeholders’ array containing objects with ‘‘name’’, ‘‘description’’, and ‘‘demographics’’ fields. The conversation is as follows:
@@
E.2 Stakeholder Reflections Prompt
@@
Appendix F Generating Stakeholder Questions
@@
Appendix G Institutional Review Board
All studies were approved by the Institutional Review Board (IRB) at anonymous institution. We obtained informed consent from all participants and participants were compensated for their time.